OpenStudy (anonymous):

a[1] = 2 and a[n] =5a[n-1] for n ≥2 (recursion formula) Write the first four terms of each sequence??

OpenStudy (amistre64):

An = 5(An-1) An = 5^2(An-2) An = 5^3(An-3) An = 5^n(A0) then right?

OpenStudy (amistre64):

or is it: An = 5^n+1 (A1).... lets make a table and determine that.. A1 = 2 A2 = 10 A3 = 50 A4 = 250 A5 = 6250 right?

OpenStudy (amistre64):

A5 = 2.5^4 = 1250

OpenStudy (amistre64):

lol.... 6250; thats a typo

OpenStudy (amistre64):

250*5 = 1250

OpenStudy (amistre64):

A4 = 2.5^3 = 250; yay!! discrete math works